Fermi National Laboratory


Accelerator Update

Friday August 6
The day shift began with store #3711 being terminated at 9:00 AM, the Tevatron (Tev) was then prepared for access to look at D1 vacuum problems. Beam continues to Switchyard 120 and MiniBoone.

At 10:12 AM problems continue with the Debuncher kicker (D:EKIK), techs have decided to change it out.

At 1:00 PM Cryo techs completed the change out of a kautzky valve in the Tev. Cool down has started.

Debuncher kicker work was completed by 3:00 PM, we then returned to Anti Proton Source (Pbar) stacking.

The Tev was back on by 6:30 PM. Studies in the Tev have begun.

Tev shot setup began at 11:33 PM.

Saturday August 7
The Midnight shift began with the Tev in shot setup, and with beam to MiniBoone and Mtest.

At 3:20 AM Main Injector (MI) RF8 supply tripped off and continued to trip when reset. The supply was removed from the loop.

Store #3715 was colliding by 3:27 AM with an initial average luminosity of 59.77 E30.

At 3:31 AM Pbar returns to stacking.

At 11:31 PM we began experiencing controls problem in switchyard, the problem was traced to a bad crate power supply, supply was replaced.

Sunday August 8
The Midnight shift began with store #3715 colliding, stacking, and with beam to MiniBoone and switchyard.

At 6:54 AM there was a vacuum burst in the Booster, when vacuum remained bad experts were called in to investigate.

At 10:46 AM an access into Booster revealed a leaking interface.

Beam returned to Booster at 4:00 PM, after repair and vacuum pump down was complete.

Store #3715 was terminated at 5:58 PM.

Tevatron shot setup began at 6:13 PM, with store #3717 colliding at 8:47 PM with an initial average luminosity of 76.92 E30

Pbar stacking and beam to MCenter and MTest has resumed.

Monday August 9
The Midnight shift began with store #3717 colliding, stacking, and with beam to MiniBoone, and switchyard 120.

At 1:13 AM we began shot setup to the Recycler Ring, this was completed by 1:35 AM.

Stacking has resumed.

At 4:06 AM, Booster was running with 2 RF (BRF4 and BRF5) stations down.

Today's Plans
Continue Pbar stacking, with some Pbar studies this evening. Keep the current Tev store until tomorrow morning.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/9/2004   email Fermilab